OpenStudy (anonymous):

a+2b+5 and a2+4b2=13, what is the value of ab=?[show each steps]

OpenStudy (alexwee123):

just one questions is it a^2 or 2(a) for the second equation?

OpenStudy (anonymous):

no simple algebra

OpenStudy (alexwee123):

so it's 2a?

OpenStudy (alexwee123):

nm lets just assume its a^2...

OpenStudy (alexwee123):

so what do you think you can do to solve a system of equations? @heartoftheocean

OpenStudy (anonymous):

just clarify\[a+2b=5\]\[a^2+4b^2=13\]what is the value of \(ab\)

OpenStudy (anonymous):

(a+2b)2 = a2+4b2 + 4ab

OpenStudy (anonymous):

i would say start by squaring both sides of \(a+2b=5\)